# Beyond Calibration: Assessing the Probabilistic Fit of Neural Regressors via Conditional Congruence

### v2

@@ Abstract
While significant progress has been made in specifying neural networks capable of representing uncertainty, deep networks still often suffer from overconfidence and misaligned predictive distributions. Existing approaches for addressing this misalignment are primarily developed under the framework of calibration, with common metrics such as Expected Calibration Error (ECE). However, calibration can only provide a strictly marginal assessment of probabilistic alignment. Consequently, calibration metrics such as ECE are distribution-wise measures and cannot diagnose the point-wise reliability of individual inputs, which is important for real-world decision-making. We propose a stronger condition, which we term conditional congruence, for assessing probabilistic fit. We also introduce a metric, Conditional Congruence Error (CCE), that uses conditional kernel mean embeddings to estimate the distance, at any point, between the learned predictive distribution and the empirical, conditional distribution in a dataset. We show that using CCE to measure congruence 1) accurately quantifies misalignment between distributions when the data generating process is known, 2) effectively scales to real-world, high dimensional image regression tasks, and 3) can be used to gauge model reliability on unseen instances.
@@ 1 INTRODUCTION
As more machine learning technologies are deployed into the real world, increasing attention is being paid to the ability of these systems to accurately represent their uncertainty [1]. Faithful uncertainty quantification improves the safety, robustness, and reliability of an ML system in the face of imperfect information about the world [2, 3] and has applications to out-of-distribution detection [4, 5, 6], active learning [7, 8], and reinforcement learning [2, 9].
@@
Figure 1: An illustration of the shortcomings of using ECE to measure probabilistic fit in regression problems. We simulate data with the following data-generating process: $X\sim\mathcal{N}(0,1)$, $Y|X\sim\mathcal{N}(\alpha X,1)$, $\alpha=3$, and introduce a poorly specified model, $f$, that ignores the conditional relationship between $X$ and $Y$, but has the same marginal: $f(x)\sim\mathcal{N}(0,1+\alpha^{2})$ (see Appendix A.1 for derivation). (a) We plot samples drawn from the true data-generating distribution, $Y|X$, along with samples from the model $f$; clearly $f$ is misspecified. (b) We marginalize out $X$ from $Y|X$ and plot the histogram of the marginal of $Y$, next to the PDF of the model, $f$, showing their statistical equivalence. (c) We compute ECE and plot the associated “reliability” diagram, achieving near-perfect calibration under ECE. (d) Only an input-aware measure of congruence, like CCE, correctly recovers the discrepancy between our model and ground-truth distributions and assesses probabilistic fit at each $x_{i}$.
While a variety of methods exist for building uncertainty-aware ML systems, in recent years training neural networks to fit high dimensional probability distributions via maximum likelihood has emerged as an effective strategy [10]. For example, a network trained to minimize cross entropy also maximizes the likelihood of a categorical distribution over the labels. In the regression setting, one can parameterize the mean and variance of a Gaussian distribution [11, 12, 13] as the outputs of a neural network, $\left[\hat{\mu}_{i},\hat{\sigma}_{i}^{2}\right]^{T}=\mathbf{f}_{\mathbf{\Theta}}(\textbf{x}_{i})$ and maximize the likelihood of the regression targets. When modeling discrete outputs, a similar approach can be employed using a Poisson [14], Negative Binomial [15], or Double Poisson [16] distribution. While these models generally achieve good mean accuracy, they have been shown to produce misaligned predictive distributions and in many cases exhibit overconfident behavior [17, 18].
@@
Such probabilistic misalignment is frequently studied through the framework of calibration, which broadly means when a model assigns a probability $p$ to a certain event, in the long run, that event should occur $p\%$ of the time. Historically, the notion of calibration was studied in the context of classification [19], and formalized under the Expected Calibration Error (ECE) metric [20, 21]. More recently, [22] propose an ECE analog for regression problems where empirical quantiles of test data are compared to a model’s predicted CDF. Regression-based calibration has received less attention than classification, despite having many relevant applications such as crowd counting [23, 24], traffic flow estimation [25, 26, 27], agricultural yield prediction [28], inventory management [29], and cellular bacterial prediction [30]
@@
While ECE is widely used to measure calibration, it has significant shortcomings. Fundamentally, it is a distribution-wise measure, which reduces calibration to a single, aggregated statistic over an entire dataset. In doing so, it can only provide a diagnostic on the marginal distribution of the regression targets, $P_{Y}$, and is unable to measure the alignment of the conditional distribution, $P_{Y|X}$. In fact, a regressor that ignores the input and predicts the marginal distribution of the targets for every instance will be considered calibrated under the paradigm introduced by [22]. See Figure 1 for an example. Moreover, this aggregation sacrifices meaningful signal at the level of each instance; it is unable to assess how well a model communicates uncertainty about a single prediction, $f(x_{i})$. Previous work has pointed out “being able to assess the reliability of a probability score for each instance is much more powerful than assigning an aggregate reliability score” [31]. Consider for example, when machine learning is applied to healthcare domains; a patient is likely much more interested in a model’s reliability for her specific case than the average reliability of the system across a test dataset [32].
To address these issues, we propose to measure the probabilistic fit of machine learning regression models with a stronger, quantifiable condition, called congruence. In contrast to calibration, which is an aggregated, distribution-wise measure of the marginal $P_{Y}$, congruence assesses the point-wise uncertainty of the conditional distribution at each specific point, $P_{Y|X=x_{i}}$. To measure congruence, we propose a novel metric called Conditional Congruence Error (CCE). Rather than rely on quantile- or coverage-based measurements, we leverage recent theoretical advances in conditional kernel mean embeddings [33] to compute conditional congruence by estimating the discrepancy between a model’s learned, predictive distribution and the empirical, conditional distribution in a dataset. Our approach does not require distributional assumptions about the data or the evaluation of a likelihood. We show that using CCE to measure point-wise congruence 1) recovers the true discrepancy between the predictive and ground truth distributions when the data generating process is known, 2) scales to high dimensional problems, such as image regression, and 3) can be used to assess the uncertainty of unseen instances.
@@ 2.1 Notation
Suppose we are given an unknown data-generating distribution $(X,Y)$, where $Y$ has a conditional relationship with $X$, $Y|X$. Let $X$ take values $x\in\mathcal{X}$ and let $Y$ take values $y\in\mathcal{Y}$. We denote $\mathcal{P}_{\mathcal{Y}}$ to be the space of all possible distributions over $\mathcal{Y}$. Let $P_{Y|X}:\mathcal{X}\rightarrow\mathcal{P}_{\mathcal{Y}}$ be nature’s mapping implied by $(X,Y)$ that takes conditioning values in $\mathcal{X}$ and outputs distributions over $\mathcal{Y}$. By $P_{Y|X=x}\in\mathcal{P}_{\mathcal{Y}}$, we refer to the specific, conditional distribution obtained over $\mathcal{Y}$ given a realization of $X$, $x\in\mathcal{X}$. $P_{X}$ and $P_{Y}$ denote the marginal distributions of $X$ and $Y$. We assume access to a probabilistic model $f:\mathcal{X}\rightarrow\mathcal{P}_{\mathcal{Y}}$, as well as a set $\mathcal{S}$ of $n$ draws from $(X,Y)$, $(x_{i},y_{i})_{i=1}^{n}$, that have been withheld from model training.
@@ 2.2 Assessing Probabilistic Fit
As discussed above, calibration is not a sufficient condition for a well-fit probabilistic model. Rather than use calibration to evaluate a model, we propose a point-wise condition that more fully captures the desiderata of supervised learning — a model that closely approximates (or is congruent with) the true conditional distribution for any input-output pair:
@@
Quantifying conditional congruence is straightforward — for all $x\in\mathcal{X}$, we compute the distance between the model’s learned conditional distribution $f(x)$ and the ground-truth $P_{Y|X=x}$. Adopting the terminology of [34] (which proposes a distance-based definition of marginal calibration in the binary classification setting), given a model $f$, we quantify its “distance from conditional congruence” ($D_{C}$) as follows:
@@
where $\rho:\mathcal{P}_{\mathcal{Y}}\times\mathcal{P}_{\mathcal{Y}}\rightarrow\mathbb{R}$ measures distance between two probability distributions over $\mathcal{Y}$. Clearly, $D_{C}(f,P_{Y|X})=0$ if and only if $f$ is conditionally, or point-wise, congruent.
@@ 3.1 Maximum Conditional Mean Discrepancy (MCMD)
In machine learning problems that model with uncertainty, we are only given a finite number of draws from $(X,Y)$ to evaluate the quality of a predictive distribution. This introduces great difficulty in estimating $D_{C}(f,P_{Y|X})$, because we require multiple examples with the exact same conditioning input $x$ to approximate $P_{Y|X=x}$ with any degree of confidence [34]. In most domains, the probability of this occurrence across multiple inputs is essentially zero.
@@
[…]
The MMD was extended to the conditional probability setting by [35], resulting in the Maximum Conditional Mean Discrepancy (MCMD). Let $P_{Y|X}$ and $P_{Y^{\prime}|X^{\prime}}$ be two mappings from $\mathcal{X}$ to $\mathcal{P}_{\mathcal{Y}}$ implied by the (possibly equivalent) joint distributions $(X,Y)$ and $(X^{\prime},Y^{\prime})$. Then, given conditional kernel mean embeddings (defined for each $x\in\mathcal{X}$), we have $\text{MCMD}(P_{Y|X},P_{Y^{\prime}|X^{\prime}},\cdot)=\|\mathcal{U}_{Y|X}(\cdot)-\mathcal{U}_{{Y^{\prime}|X^{\prime}}}(\cdot)\|_{\mathcal{H}_{\mathcal{Y}}}$. Recently, [33] derived a closed-form estimate of MCMD, which we provide in Equation 2.
@@ 3.1.1 Computing MCMD
Given a set of draws $\mathcal{S}=(x_{i},y_{i})_{i=1}^{n}\sim(X,Y)$ and $\mathcal{S}^{\prime}=(x^{\prime}_{i},y^{\prime}_{i})_{i=1}^{m}\sim(X^{\prime},Y^{\prime})$, we have the following closed-form, sample-based estimate from [33] for the squared MCMD at a given conditioning value $(\cdot)\in\mathcal{X}$:
@@
where we first choose input and output kernels $k_{\mathcal{X}}$ and $k_{\mathcal{Y}}$, yielding $\textbf{K}_{Y}^{(ij)}=k_{\mathcal{Y}}(y_{i},y_{j})$, $\textbf{K}_{Y^{\prime}}^{(ij)}=k_{\mathcal{Y}}(y^{\prime}_{i},y^{\prime}_{j})$, $\textbf{K}_{YY^{\prime}}^{(ij)}=k_{\mathcal{Y}}(y_{i},y^{\prime}_{j})$, $\textbf{K}_{X}^{(ij)}=k_{\mathcal{X}}(x_{i},x_{j})$, $\textbf{K}_{X^{\prime}}^{(ij)}=k_{\mathcal{X}}(x^{\prime}_{i},x^{\prime}_{j})$, $\textbf{k}_{X}(\cdot)$ = $[k_{\mathcal{X}}(x_{1},\cdot),...,k_{\mathcal{X}}(x_{m},\cdot)]^{T}$, $\textbf{k}_{X^{\prime}}(\cdot)$ = $[k_{\mathcal{X}}(x^{\prime}_{1},\cdot),...,k_{\mathcal{X}}(x^{\prime}_{m},\cdot)]^{T}$, $\textbf{W}_{X}=(\textbf{K}_{X}+n\lambda\textbf{I}_{n})^{-1}$ and $\textbf{W}_{X^{\prime}}=(\textbf{K}_{X^{\prime}}+m\lambda^{\prime}\textbf{I}_{m})^{-1}$
Both $\lambda$ and $\lambda^{\prime}$ are regularizing hyperparameters that we set to 0.1 throughout this paper. We study the effect of $\lambda$ in Appendix C.3.
@@ 3.1.2 Theoretical Guarantees of MCMD
[33] provide a handful of theoretical guarantees associated with the MCMD. The most relevant to our work (Theorem 5.2 in the original paper) is restated below, with adjustments to match our notation:
@@
This is a useful theorem, and as stated by [33], does not have restrictive assumptions. We note that the guarantees are given in terms of $\text{MCMD}(P_{Y|X},P_{Y^{\prime}|X^{\prime}},\cdot)$ and not the estimate $\widehat{\text{MCMD}^{2}}$. However, if we also assume bounded kernels (once again, not restrictive), [33] provide several desirable convergence properties with their estimate (see Theorems 4.4, 4.5 in their paper) that give us confidence in employing the MCMD as a measure of congruence.
If a model $f$ yields samples that achieve near-zero MCMD when compared to the ground-truth samples, these theoretical results ensure that the model has learned a conditional probability distribution with an almost-sure equality to the ground-truth data-generating distribution. In fact, this allows us to use the MCMD to directly estimate $D_{C}(f,P_{Y|X})$, as detailed in the next section.
Figure 2: Visualization of the probabilistic fit of 4 deep neural networks on a synthetic dataset with discrete regression targets. (Row 1) A plot of the mean and centered 95% credible interval for each model, along with the ECE and NLL. (Row 2) CCE between the ground-truth and predictive distributions, along with the mean value ($\overline{\mbox{CCE}}$) across the evaluation grid.
3.2 Conditional Congruence Error
While the MCMD is a powerful theoretical tool, applying it to high-dimensional joint distributions common to deep learning introduces two main challenges: 1) representation of the inputs in a space that lends itself to kernelized similarities; and 2) a method for defining (and obtaining samples from) $(X^{\prime},Y^{\prime})$ given only a test set $S=(x_{i},y_{i})_{i=1}^{n}$ and a model $f$. We specify a process that addresses both challenges, thus yielding a metric, the Conditional Congruence Error (CCE), that provides a holistic look at a model’s probabilistic fit across the input space.
To compute the CCE, we first specify an encoding function $\varphi:\mathcal{X}\rightarrow\mathbb{R}^{d}$ that can map each complex input (images, text, etc.) to a fixed-length vector in a semantic embedding space, similar to [36] and [37]. We then embed each input in our test set and treat $\mathcal{S}_{\varphi}=(\varphi(x_{i}),y_{i})_{i=1}^{n}$ as our sample from nature’s distribution $(\varphi(X),Y)$. To obtain $\mathcal{S}^{\prime}_{\varphi}=(\varphi(x^{\prime}_{i}),y^{\prime}_{i})_{i=1}^{m}\sim(\varphi(X^{\prime}),Y^{\prime})$, we draw $\ell$ Monte Carlo samples from each predictive distribution $f(x_{i})$. This forms a collection of $m=n\ell$ draws $\{(\varphi(x_{i}),y_{ik})_{k=1}^{\ell}\}_{i=1}^{n}$ from the model’s learned conditional distributions which we can then compare with the ground-truth draws. Armed with these samples, we can now define CCE:

### v3

# Assessing the Probabilistic Fit of Neural Regressors via Conditional Congruence

@@ Abstract
While significant progress has been made in specifying neural networks capable of representing uncertainty, deep networks still often suffer from overconfidence and misaligned predictive distributions. Existing approaches for measuring this misalignment are primarily developed under the framework of calibration, with common metrics such as Expected Calibration Error (ECE). However, calibration can only provide a strictly marginal assessment of probabilistic alignment. Consequently, calibration metrics such as ECE are distribution-wise measures and cannot diagnose the point-wise reliability of individual inputs, which is important for real-world decision-making. We propose a stronger condition, which we term conditional congruence, for assessing probabilistic fit. We also introduce a metric, Conditional Congruence Error (CCE), that uses conditional kernel mean embeddings to estimate the distance, at any point, between the learned predictive distribution and the empirical, conditional distribution in a dataset. We perform several high dimensional regression tasks and show that CCE exhibits four critical properties: correctness, monotonicity, reliability, and robustness.
[…]
2886
@@ 1 Introduction
As more machine learning (ML) technologies are deployed into the real world, increasing attention is being paid to the ability of these systems to accurately represent their uncertainty [1]. Faithful uncertainty quantification (UQ) improves the safety, robustness, and reliability of an ML system in the face of imperfect information about the world [47, 23] and has applications to out-of-distribution detection [2, 25], active learning [41], and reinforcement learning [47, 22]. Accordingly, it is crucial not only to measure a model’s uncertainty, but to assess the quality of the UQ measurement.
@@
Figure 1: An illustration of the shortcomings of using ECE to measure probabilistic fit in regression problems. We simulate a model, $f$, that ignores the conditional relationship between $X$ and $Y$, but whose predictions match the ground-truth marginal distribution of $Y$ (see Appendix B.1). (a) We plot samples drawn from the true data-generating distribution, $Y|X$, along with samples from the model $f$; clearly $f$ is misspecified. (b) We marginalize out $X$ from $Y|X$ and plot the histogram of the marginal of $Y$, next to the PDF of the model, $f$, showing their statistical equivalence. (c) We compute ECE and plot the associated “reliability” diagram, achieving near-perfect calibration under ECE. (d) CCE correctly recovers the discrepancy between our model and ground-truth distributions and assesses probabilistic fit at each $x_{i}$.
While diagnosing the quality of UQ estimates in classification tasks is well-studied [16, 30], historically the regression setting has received less attention despite many relevant applications. A popular approach to capture uncertainty is to train a deep neural network (DNN) to fit a probability distribution over the labels, conditioned on the input. For example, prior work parameterizes the mean and variance of a Gaussian distribution [34, 3, 40, 42] as the outputs of a neural network and maximizes the likelihood of the regression targets. When modeling count data, a similar approach can be employed using a Poisson [14], Negative Binomial [45], or Double Poisson [46] distribution. To improve accuracy and UQ quality, multiple independently trained networks can be combined as deep ensembles [28]. While probabilistic DNNs generally achieve good mean accuracy, they have been shown in many cases to produce misaligned or overconfident predictive distributions [16, 30]. Therefore, accurate diagnostics of UQ quality are of paramount importance.
@@
Such probabilistic misalignment is frequently studied through the framework of calibration, which broadly requires that when a model assigns a probability $p$ to a certain event, in the long run, that event should occur $(100p)\%$ of the time. Historically, calibration was studied in the context of classification [9], and formalized under the Expected Calibration Error (ECE) metric [33, 35]. More recently, Kuleshov et al. [26] proposed an ECE analog for regression problems where empirical quantiles of test data are compared against a model’s predicted CDFs.
@@
While ECE is widely used to measure calibration on regression tasks, it has significant shortcomings. Fundamentally, it is a distribution-wise measure, which can only provide diagnostics on the marginal predictions of a model while being unable to quantify the alignment of its conditional distributions. In fact, a regressor that ignores the input and predicts the marginal distribution of the targets for every instance will be considered calibrated under the paradigm introduced by Kuleshov et al. [26]. See Figure 1 for an example. Moreover, ECE sacrifices meaningful signal at the level of each instance; it is unable to make point-wise assessments of a model’s uncertainty for a single prediction. Previous work has pointed out “being able to assess the reliability of a probability score for each instance is much more powerful than assigning an aggregate reliability score” [27]. Consider, for example, when machine learning is applied to healthcare domains; a patient is likely much more interested in a model’s reliability for her specific case than its average reliability across a test set [19].
An alternative to ECE is to employ a proper scoring rule, such as the Negative Log Likelihood (NLL), to measure the quality of a predictive model. NLL is attractive because it can be used both distribution-wise and point-wise to assess probabilistic fit. However, NLL has several drawbacks that make it difficult or inappropriate to use in practice. First, a well-known artifact of likelihood-based estimators is that they are sensitive to outliers [32]. Second, NLL cannot be used to compare models with different distributional assumptions and therefore different likelihoods [7]. Third, NLL has no known lower bound, which makes it difficult to interpret just how close to perfectly-fit a given predictive distribution is [5]. Lastly, while NLL can be used point-wise to assess probabilistic fit, it requires access to a label at the diagnostic point to do so. This restricts the usefulness of NLL in settings where we need to estimate the reliability of a model’s forecast on a novel input.
To address the limitations of both calibration- and likelihood-based diagnostics, we propose to assess the probabilistic fit of regression models with a distance-based condition, which we call congruence. In contrast to calibration, which is an aggregated, distribution-wise measure of a model’s marginal predictions, congruence considers the uncertainty quality of each point-wise conditional distribution. To quantify congruence, we propose a novel metric called Conditional Congruence Error (CCE). This metric leverages recent theoretical advances in conditional kernel mean embeddings [37] to estimate the discrepancy between a model’s learned, predictive distribution and the empirical, conditional distribution in a dataset. Our method can be used to evaluate point-wise conditional misalignment, is able to compare models with different distributional assumptions, and is lower bounded by 0. In our experiments we show that CCE displays four critical properties of a desirable UQ diagnostic, including correctness, monotonicity, reliability, and robustness. Finally, we include case studies describing how CCE can be used to judge the point-wise reliability of predictions on unseen instances without access to labels.
@@ 2.1 Notation
We consider an unknown data-generating distribution $(X,Y)$, where $Y$ has a conditional relationship with $X$, i.e. $Y|X$. $X$ takes values $x\in\mathcal{X}$, and $Y$ takes values $y\in\mathcal{Y}$. We use $\mathcal{P}_{\mathcal{Y}}$ to denote the space of all possible distributions over $\mathcal{Y}$. Let $P_{Y|X}:\mathcal{X}\rightarrow\mathcal{P}_{\mathcal{Y}}$ be nature’s mapping that takes conditioning values in $\mathcal{X}$ and outputs distributions over $\mathcal{Y}$. By $P_{Y|X=x}$, we refer to the specific distribution obtained over $\mathcal{Y}$ given a realization of $X$, $x\in\mathcal{X}$. $P_{X}$ and $P_{Y}$ denote the marginal distributions of $X$ and $Y$. We assume access to a probabilistic model $f:\mathcal{X}\rightarrow\mathcal{P}_{\mathcal{Y}}$, as well as a set $\mathcal{S}$ of $n$ samples from $(X,Y)$, $(x_{i},y_{i})_{i=1}^{n}$, that have been withheld from model training.
@@ 2.2 Assessing probabilistic fit
Since calibration is not a sufficient condition for a well-fit probabilistic model, we propose to evaluate with a point-wise condition that more fully captures how well a model approximates (or is congruent with) the true conditional distribution for any input-output pair:
@@
Quantifying conditional congruence is straightforward — for all $x\in\mathcal{X}$, we compute the distance between the model’s learned conditional distribution $f(x)$ and the ground-truth $P_{Y|X=x}$. Adopting the terminology of Błasiok et al. [5] (which proposes a distance-based definition of marginal calibration in the binary classification setting), given a model $f$, we quantify its “distance from conditional congruence” ($D_{C}$) as follows:
@@
where $\rho:\mathcal{P}_{\mathcal{Y}}\times\mathcal{P}_{\mathcal{Y}}\rightarrow\mathbb{R}$ measures the distance between two probability distributions over $\mathcal{Y}$. Clearly, $D_{C}(f,P_{Y|X})=0$ if and only if $f$ is conditionally, or point-wise, congruent.
@@ 3.1 Maximum Conditional Mean Discrepancy (MCMD)
When assessing a model, we are only given a finite number of draws from $(X,Y)$ to evaluate the quality of its predictive distributions. This introduces great difficulty in estimating $D_{C}(f,P_{Y|X})$, because we require several examples with the exact same input $x$ to reasonably approximate $P_{Y|X=x}$ [5]. In most domains, the probability of this occurrence across multiple inputs is essentially zero.
@@
Kernel mean embeddings [31] extend the idea of applying a higher-dimensional feature map, $\phi$, to the space of probability distributions, $\mathcal{P}_{X}$. Let $P\in\mathcal{P}_{\mathcal{X}}$ be a probability distribution over $\mathcal{X}$. Then the kernel mean embedding is defined as $\phi(P)=\mathcal{U}_{P}\triangleq\int_{\mathcal{X}}k(\textbf{x},\cdot)dP(\textbf{x})$. In this formulation, $P$ is represented as $\mathcal{U}_{P}\in\mathcal{H}$, an element of an infinite-dimensional reproducing kernel Hilbert space (RKHS) [31]. One application of this technique is the maximum mean discrepancy (MMD), which estimates the distance between two distributions, $P\in\mathcal{P}_{\mathcal{X}}$ and $Q\in\mathcal{P}_{\mathcal{X}}$, by computing the magnitude of the difference between their respective mean embeddings: $\text{MMD}(P,Q)=\left\lVert\mathcal{U}_{P}-\mathcal{U}_{Q}\right\rVert_{\mathcal{H}}$. Using the kernel trick we can estimate this distance (which is an inner product) from data points without having to explicitly evaluate the mapping $\phi(x)$.
The MMD was extended to the conditional probability setting by Muandet et al. [31], resulting in the Maximum Conditional Mean Discrepancy (MCMD). Let $P_{Y|X}$ and $P_{Y^{\prime}|X^{\prime}}$ be two mappings from $\mathcal{X}$ to $\mathcal{P}_{\mathcal{Y}}$ implied by the (possibly equivalent) joint distributions $(X,Y)$ and $(X^{\prime},Y^{\prime})$. Then, given conditional kernel mean embeddings (defined for each $x\in\mathcal{X}$), we have $\text{MCMD}(P_{Y|X},P_{Y^{\prime}|X^{\prime}},\cdot)=\|\mathcal{U}_{Y|X}(\cdot)-\mathcal{U}_{{Y^{\prime}|X^{\prime}}}(\cdot)\|_{\mathcal{H}_{\mathcal{Y}}}$. Recently, Park and Muandet [37] derived a closed-form estimate of MCMD, which we provide in Equation 2.
@@ 3.1.1 Computing MCMD
Given a set of draws $\mathcal{S}=(x_{i},y_{i})_{i=1}^{n}\sim(X,Y)$ and $\mathcal{S}^{\prime}=(x^{\prime}_{i},y^{\prime}_{i})_{i=1}^{m}\sim(X^{\prime},Y^{\prime})$, we have the following sample-based estimate for the squared MCMD at a given conditioning value $(\cdot)\in\mathcal{X}$:
@@
where we first choose input and output kernels $k_{\mathcal{X}}$ and $k_{\mathcal{Y}}$, yielding $\textbf{K}_{Y}^{(ij)}=k_{\mathcal{Y}}(y_{i},y_{j})$, $\textbf{K}_{Y^{\prime}}^{(ij)}=k_{\mathcal{Y}}(y^{\prime}_{i},y^{\prime}_{j})$, $\textbf{K}_{YY^{\prime}}^{(ij)}=k_{\mathcal{Y}}(y_{i},y^{\prime}_{j})$, $\textbf{K}_{X}^{(ij)}=k_{\mathcal{X}}(x_{i},x_{j})$, $\textbf{K}_{X^{\prime}}^{(ij)}=k_{\mathcal{X}}(x^{\prime}_{i},x^{\prime}_{j})$, $\textbf{k}_{X}(\cdot)$ = $k_{\mathcal{X}}(x_{1},\cdot),...,k_{\mathcal{X}}(x_{n},\cdot)]^{T}$, $\textbf{k}_{X^{\prime}}(\cdot)$ = $k_{\mathcal{X}}(x^{\prime}_{1},\cdot),...,k_{\mathcal{X}}(x^{\prime}_{m},\cdot)]^{T}$, $\textbf{W}_{X}=(\textbf{K}_{X}+n\lambda\textbf{I}_{n})^{-1}$ and $\textbf{W}_{X^{\prime}}=(\textbf{K}_{X^{\prime}}+m\lambda^{\prime}\textbf{I}_{m})^{-1}$
Both $\lambda$ and $\lambda^{\prime}$ are regularizers that we set to 0.1 throughout this paper. We study the effect of $\lambda$ in Appendix A.5.
@@ 3.1.2 Theoretical guarantees of MCMD
Park and Muandet [37] provide a handful of theoretical guarantees associated with the MCMD. We restate the most relevant to our work (Theorem 5.2 in the original paper), adjusted to match our notation:
@@
This is a useful theorem, without restrictive assumptions. We note that the guarantees are given in terms of $\text{MCMD}(P_{Y|X},P_{Y^{\prime}|X^{\prime}},\cdot)$ and not the estimate $\widehat{\text{MCMD}^{2}}$. However, if we also assume bounded kernels, we obtain several desirable convergence properties of this estimate (see Theorems 4.4, 4.5 in [37]) that give us confidence in employing the MCMD as a measure of congruence.
If a model $f$ yields samples that achieve near-zero MCMD when compared to ground-truth data, these theoretical results suggest that $f$ has learned a conditional distribution with an almost-sure equality to nature’s data-generating distribution. This allows us to use the MCMD to directly estimate $D_{C}(f,P_{Y|X})$, as detailed in the next section.
3.2 Conditional Congruence Error (CCE)
While MCMD is a powerful theoretical tool, applying it to the high-dimensional joint distributions common in modern ML introduces two challenges: 1) representation of the inputs in a space that lends itself to kernelized similarities; and 2) defining (and obtaining samples from) $(X^{\prime},Y^{\prime})$ given only a test set $S=(x_{i},y_{i})_{i=1}^{n}$ and a model $f$. We specify a process that addresses both challenges, thus yielding a metric, the Conditional Congruence Error (CCE), that provides a holistic look at a model’s probabilistic fit across the input space.
To compute CCE, we first specify an encoding function $\varphi:\mathcal{X}\rightarrow\mathbb{R}^{d}$ that can map each complex input (images, text, etc.) to a fixed-length vector in a semantic embedding space [39]. We then embed each input in our test set and treat $\mathcal{S}_{\varphi}=(\varphi(x_{i}),y_{i})_{i=1}^{n}$ as our sample from nature’s distribution $(\varphi(X),Y)$. To obtain $\mathcal{S}^{\prime}_{\varphi}=(\varphi(x^{\prime}_{i}),y^{\prime}_{i})_{i=1}^{m}\sim(\varphi(X^{\prime}),Y^{\prime})$, we draw $\ell$ Monte Carlo samples from each predictive distribution $f(x_{i})$. This forms a collection of $m=n\ell$ draws $\{(\varphi(x_{i}),y_{ik})_{k=1}^{\ell}\}_{i=1}^{n}$ from the model’s learned conditional distributions which we can then compare with the ground-truth draws. Armed with these samples, we can now define CCE: